what values of A make the relation (A,1), (2,3), (4,5) a function?

Answer by checkley77(12844)

slope=(5-3)/(4-2)=2/2=1
(3-1)/(2-x)=1
2=1(2-x)
2=2-x
-x=2-2
x=0 ans. all three points lie on the same straight line.
(0,1), (2,3), (4,5)
Y=X+1 is the equation of this line.